OpenStudy (anonymous):

Using the equation, 4Fe + 3O2 yields to 2Fe2O3, if 111.6 g of iron react, what mass of oxygen was consumed?

OpenStudy (aaronq):

Convert everything to moles (in stoichiometry, it needs to be in moles) Then use the stoichiometric coefficients to build a ratio in this manner: e.g. for a general reaction: aA + bB \(\rightleftharpoons\) cC where upper case are the species, and lower case are the coefficients, \(\dfrac{n_A}{a}=\dfrac{n_B}{b}=\dfrac{n_C}{c}\) from here, you can solve for moles of oxygen by plugging in the moles of iron. Then convert moles to grams.

OpenStudy (anonymous):

thank you so much for the help!

OpenStudy (aaronq):

no problem !

OpenStudy (anonymous):

@aaronq would you happen to know how to do this problem? Write the proper word equation to express the following chemical reaction: 3Li (s) + AuCl3 (aq) yields to 3LiCl (aq) + Au (s).

OpenStudy (aaronq):

you just gotta turn those formulas into words. the notation in brackets (s) and (aq) imply their states. (s) means solid, as in Li(s) lithium metal and AuCl3 would be Gold(III)chloride

OpenStudy (aaronq):

aqueous gold(III)chloride***